# 1D Transition Metal Oxide Chains as a Challenging Model for Ab Initio Calculations

### v1

@@ Abstract
Providing highly simplified models of strongly correlated electronic systems that challenge ab initio calculations can serve as a valuable testing ground to improve these methods. In this study, we present a comprehensive study of the structural, magnetic, and electronic properties of one-dimensional transition metal mono-oxide chains (VO, CrO, MnO, FeO, CoO, and NiO) using density functional theory (DFT), DFT+$U$, and coupled-cluster singles and doubles (CCSD) calculations. The Hubbard $U$ parameter for DFT+$U$ is determined using the linear response theory. In all systems studied except MnO, the presence of multiple local minima —primarily due to the electronic degrees of freedom associated with the d-orbitals— leads to significant challenges for DFT, DFT+U, and Hartree-Fock methods in finding the global minimum in ab initio calculations. Our results indicate that the antiferromagnetic (AFM) state is energetically favored for all chains, except CrO, when using DFT+$U$ and PBE. We analyze the electronic band structures and find that while the PBE approximation often predicts metallic or half-metallic ground states for the ferromagnetic (FM) state, DFT+$U$ approach successfully opens band gaps, correctly predicting insulating behavior in all cases. Furthermore, we compared the energy differences between the AFM and FM states using DFT+$U$ and CCSD for CrO, MnO, FeO, CoO and NiO. Our findings indicate that CCSD predicts larger energy differences in some cases compared to DFT+$U$, suggesting that the Hubbard $U$ parameter obtained through linear response theory may be overestimated when used to calculate energy differences between different magnetic states. For CrO, CCSD predicts an AFM ground state, in contrast to the predictions from DFT+$U$ and PBE methods.
@@ I Introduction
A major challenge in DFT+$U$ is determining the Hubbard $U$ parameter accurately to ensure consistency with experimental results. Recently, we demonstrated that using linear response theory to estimate $U$  Cococcioni and de Gironcoli (2005) can lead to an underestimation of Heisenberg exchange interactions, which, in turn, results in underestimated transition temperatures for certain antiferromagnetic materials, particularly 3d transition metal oxides (TMOs) Mosleh and Alaei (2023).
@@
While DFT+$U$ offers improvements over standard DFT, it is essential to assess how effectively this approach —and other ab initio methods— captures the intricate magnetic properties of TMOs. These properties are closely tied to the choice of exchange-correlation approximations. Comparing DFT+$U$ results with more sophisticated methods is a key step toward refining its accuracy. However, modeling bulk TMOs is computationally demanding. A practical alternative is to study simpler systems, such as one-dimensional (1D) chains, which serve as effective testing ground for evaluating DFT+$U$ against more advanced methods.
Using 1D chains as models in ab initio studies is a well-established approach for developing and testing quantum chemistry methods. Hydrogen chains, for example, have been extensively studied in quantum chemistry and materials science Motta et al. (2020, 2017); Stella et al. (2011); Umari and Marzari (2009); Liu et al. (2011). In this work, we investigate one-dimensional transition metal oxide chains (1D-TMOs), VO, CrO, MnO, FeO, CoO and NiO, as model systems for benchmarking different computational methods. Although 1D-TMOs can be synthesized on surfaces Schmitt et al. (2019); Ferstl et al. (2016), we propose them as ideal models for testing and improving computational techniques. With fewer electrons than their bulk counterparts, 1D-TMOs provide a simplified yet realistic platform for assessing the performance of advanced ab initio methods.
Our study systematically evaluates a range of computational methods, including DFT, DFT+$U$, and the highly accurate but computationally expensive coupled-cluster method with single and double excitations (CCSD). We analyze the structural stability, electronic properties, and magnetic ordering of these 1D-TMOs, highlighting the strengths and limitations of each approach. Additionally, we examine critical issues such as the influence of pseudopotentials versus full-potential calculations, convergence challenges, and the effects of structural distortions on electronic behavior.
In summary, this paper provides a comprehensive analysis of ab initio methods applied to 1D-TMOs, offering insights into the challenges of modeling these systems. The paper is organized as follows: Section 2 details the computational methods, Section 3 presents our findings, including an analysis of convergence challenges and a comparison of DFT+$U$ with CCSD, and Section 4 concludes with a summary of our results.
@@ II Computational Methods
We use three computational codes for our calculations: Quantum ESPRESSO (QE) Giannozzi et al. (2009), a plane-wave pseudopotential code; PySCF Sun et al. (2018), a Python-based quantum chemistry framework; and FHI-aims Blum et al. (2009), an all-electron, full-potential electronic structure code. For density functional theory (DFT) calculations, we employ the Perdew-Burke-Ernzerhof (PBE) functional Perdew et al. (1996), a generalized gradient approximation (GGA) for the exchange-correlation energy. In DFT+$U$ calculations, we incorporate Dudarev’s formulation Dudarev et al. (1998) as implemented in QE.
@@
Our study focuses on first-row transition metal mono-oxides (TMOs, with TM = V, Cr, Mn, Fe, Co, and Ni) arranged in a one-dimensional (1D) chain structure along the x-direction (Fig.1). Each chain is investigated in two magnetic configurations: ferromagnetic (FM) and antiferromagnetic (AFM). For AFM states, we use a minimal unit cell containing two formula units (4 atoms) to account for magnetic ordering, and for the FM states, we adopt the same geometry unless explicitly stated. To minimize interactions between periodic images, we introduce a vacuum thickness of 30 atomic units (a.u.), and the Brillouin zone is sampled using a $4\times 1\times 1$ k-point mesh.
In QE, GBRV ultra-soft pseudopotentials are used Garrity et al. (2014) with a kinetic energy cutoff of 60 Ry for the plane-wave expansion. For all-electron calculations in FHI-aims, we select a tight-tier2 basis set. In PySCF calculations, we employ the Goedecker-Teter-Hutter (GTH) pseudopotential with the corresponding GTH-DZVP-MOLOPT-SR basis set (abbreviated as DZVP) VandeVondele and Hutter (2007); Gao et al. (2020), obtained from the CP2K library VandeVondele et al. (2005), to represent the valence electrons. Unrestricted CCSD calculations are carried out with initial reference states generated from unrestricted Hartree-Fock (UHF).
@@
Figure 1: The crystal structure of a transition metal monoxide (TMO) one-dimensional (1D) chain, oriented along the $x$-axis, is shown. The transition metal (TM = V, Cr, Mn, Fe, Co, Ni) and oxygen atoms are represented by blue and red circles, respectively. The arrows indicate the magnetic moments of the transition metal atoms, illustrating two magnetic configurations: FM and AFM states.
@@ III.1 Instability and Convergence Issues
With the exception of the MnO chain, which shows stable convergence, all DFT and DFT+$U$ calculations—regardless of the DFT code used (i.e., Pyscf, QE, and FHI-aims)—face significant wave function instability issues, often causing the self-consistent field (SCF) calculations to converge to an excited state instead of the ground state. Detecting such instabilities only requires comparing the total energy of the 1D-TMOs at different lattice parameters. Fig. 2(a) illustrates the variation in total energy with respect to the lattice parameter (TM-TM distance) for the CoO chain in the AFM state. Although the calculations converge successfully for all lattice parameters, some converge to unstable states, resulting in a zigzag energy plot rather than the expected smooth curve. This behavior indicates that some SCF calculations fail to reach a ground-state solution.
Projected density of states (PDOS) calculations reveal differences in d-orbital occupations between unstable and stable points, suggesting that these instabilities may arise from the limitations of ab initio methods in determining optimal electronic d-orbital occupations. Figure 3 presents the spin-down charge density profile of CoO in a (011) surface cut, showing significant differences between stable and unstable configurations. This suggests that the d-orbital degrees of freedom in one-dimensional transition metal oxides (1D-TMOs) can contribute to instability in ab initio calculations.
@@
To overcome this issue, we adopt a straightforward yet effective approach. First, we analyze the total energy profile (Fig. 2(a)) and the magnetic moments to identify the most stable solution. We then use the wave function of this stable configuration–characterized by the expected magnetic moments and lower energy–as the initial wave function for calculations at points prone to instability. This method enables us to bypass instability issues and achieve ground states across all lattice parameters in our calculations, as demonstrated in Fig. 2(b). We apply this approach to all DFT and HF calculations.
Figure 2: The Variation in the total energy versus lattice constant (TM-TM distance) for CoO using PBE method calculated by Quantum Espresso, in AFM phase :(a) Initial calculations showing instability in the wave function. (b) Final calculations after applying our trick to address the instability, demonstrating improved convergence.
Figure 3: (Colour online) The iso-surface of the spin density for CoO in the FM state: (a) The spin density is isotropic for the unstable configuration. (b) An anisotropic spin density is observed for the stable configuration. (c) Shows the difference in spin density (subtracted). A slice of the spin density centered on the Co atom in the (011) surface is shown in the right panel.
@@ III.2 Structural Properties
To study the electronic structure and magnetic properties of 1D-TMOs, we first determine their most stable atomic configurations. We perform a series of calculations for different lattice constants, considering both FM and AFM states. Figure 4 illustrates how the total energy varies with the lattice parameter (TM-TM distance) for each chain, computed using different ab initio methods.
Our results indicate that the AFM state is the most stable configuration for all chains, except CrO, when using both PBE and DFT+$U$. We obtain the optimized lattice constants, corresponding to the minimum energy values, by fitting the total energy data to the Birch-Murnaghan equation of state  Murnaghan (1944); Birch (1947). These optimized lattice constants are then used to determine other properties, such as the magnetic moments of the 1D-TMOs.
To calculate the magnetization, we analyze the projected density of states (PDOS) of these systems. Table 1 summarizes the ground-state lattice constants, magnetization, and energy parameter ($\Delta E$) for each chain. The table clearly shows that the FM state generally results in slightly larger lattice constants and higher magnetic moments. However, within PBE approach, the NiO chain is an exception, where the AFM state shows a larger lattice constant and higher magnetic moment compared to the FM state.
Additionally, the magnetization of NiO is lower than expected within PBE, which can be attributed to its electronic structure near the Fermi level, where the orbitals are half-filled. This finding highlights the necessity of applying a Hubbard $U$ correction through DFT+$U$ calculations.
Figure 4: The variation of total energy with respect to the lattice constant. The reference energy ($E_{\mathrm{ref}}$) corresponds to the minimum energy of the stable phase. The calculations were performed using the PBE and DFT+$U$ methods with Quantum ESPRESSO and the CCSD method with PySCF for the FM and AFM phases of 1D transition metal oxide chains.
Table 1: The calculated lattice parameters (TM-TM distance) and magnetic moments per ion, in the AFM and FM phases as well as the energy differences ($\Delta E$) between two magnetic phases using PBE method in the QE, FHI-aims and PySCF, DFT+$U$ method in QE and CCSD method in PySCF. The absence of results for VO in PBE(PySCF) is due to the need for basis sets larger than TZVP to achieve convergence. For VO, we use fixed U values of 4.8 eV, while for the other compounds, we employ a self-consistent U parameter determined using DFPT.
@@
In the next step, to validate the pseudo-potentials, we conducted a comparison between results obtained using the full-potential FHI-aims and pseudo-potentials from QE, employing the GGA exchange-correlation functional. The calculated parameters are presented in Table 1. The lattice parameters calculated using QE and FHI-aims show good agreement, with differences generally within 0.01 to 0.02 atomic units (a.u.), indicating the reliability of the present pseudo-potentials approach to predict the lattice constants. The Mulliken population analysis to estimate atomic spins also shows consistency between the two methods, though some differences are notable. For example, the magnetic moments of VO, CrO and MnO show small variations. The differences are more pronounced in compounds such as FeO (AFM), CoO (FM), and NiO (AFM). For the AFM state of FeO, the magnetization difference calculated by the two codes is 0.24 $\mu_{B}$, and for the FM state of CoO it is 0.98 $\mu_{B}$, and finally for the AFM state of NiO the difference is about 0.2 $\mu_{B}$.
The energy differences between the AFM and FM phases ($\Delta E$) are generally consistent between the two methods, although there are some discrepancies. For VO, CrO, MnO, CoO and NiO, the difference in $\Delta E$ between the methods is less than or equal to 0.05 eV in all cases. Only for FeO does this difference exceed 0.16 eV.
In summary, the GBRV pseudo-potentials in QE generally provide results in close agreement with the full-potential FHI-aims calculations for lattice parameters, magnetic moments, and energy differences, validating their use in predicting the properties of these transition metal monoxide chains. The slight variations observed underscore the nuanced differences between pseudo-potentials and all-electron methods, emphasizing the importance of method selection based on the specific requirements of a study.

### v2

@@ Abstract
Providing highly simplified models of strongly correlated electronic systems that challenge ab initio calculations can serve as a valuable testing ground to improve these methods. In this study, we present a comprehensive investigation of the structural, magnetic, and electronic properties of one-dimensional transition metal mono-oxide chains (VO, CrO, MnO, FeO, CoO, and NiO) using density functional theory (DFT), DFT+$U$, and coupled-cluster singles and doubles (CCSD) calculations. The Hubbard $U$ parameter for DFT+$U$ is determined using linear response theory. In all systems studied except MnO, the presence of multiple local minima—primarily due to the electronic degrees of freedom associated with the d-orbitals—leads to significant challenges for DFT, DFT+$U$, and Hartree–Fock methods in finding the global minimum in ab initio calculations. Our results indicate that the antiferromagnetic (AFM) state is energetically favored for all chains, except CrO, when using DFT+$U$ and the Perdew–Burke–Ernzerhof (PBE) functional. Analysis of the band structures shows that while PBE often predicts metallic or half-metallic FM states, DFT+$U$ opens band gaps and correctly yields insulating behavior in all cases. Furthermore, we compared the energy differences between the AFM and FM states using DFT+$U$ and CCSD for CrO, MnO, FeO, CoO, and NiO. Our findings indicate that CCSD predicts larger energy differences in some cases compared to DFT+$U$, suggesting that the Hubbard $U$ parameter obtained through linear response theory may be overestimated when used to calculate energy differences between different magnetic states. For CrO, CCSD predicts an AFM ground state, in contrast to the predictions from DFT+$U$ and PBE methods.
@@ I INTRODUCTION
A major challenge in DFT+$U$ is determining the Hubbard $U$ parameter accurately to ensure consistency with experimental results. Recently, we demonstrated that using linear response theory to estimate $U$ Cococcioni and De Gironcoli (2005) can lead to an underestimation of Heisenberg exchange interactions, which, in turn, results in underestimated transition temperatures for certain antiferromagnetic (AFM) materials, particularly 3d transition metal oxides (TMOs) Mosleh and Alaei (2023).
@@
While DFT+$U$ offers improvements over standard DFT, it is essential to assess how effectively this approach —and other ab initio methods— captures the intricate magnetic properties of TMOs. These properties are closely tied to the choice of exchange-correlation approximations. Comparing DFT+$U$ results with more sophisticated methods is a key step toward refining its accuracy. However, modeling bulk TMOs is computationally demanding. A practical alternative is to study simpler systems, such as one-dimensional (1D) chains, which serve as an effective testing ground for evaluating DFT+$U$ against more advanced methods.
Using 1D chains as models in ab initio studies is a well-established approach for developing and testing quantum chemistry methods. Hydrogen chains, for example, have been extensively studied in quantum chemistry and materials science Motta et al. (2020, 2017); Stella et al. (2011); Umari and Marzari (2009); Liu et al. (2011). In this work, we investigate one-dimensional transition metal oxide chains (1D-TMOs)—VO, CrO, MnO, FeO, CoO, and NiO—as model systems for benchmarking different computational methods. Although 1D-TMOs can be synthesized on surfaces Schmitt et al. (2019); Ferstl et al. (2016), we propose them as ideal models for testing and improving computational techniques. With fewer electrons than their bulk counterparts, 1D-TMOs provide a simplified yet realistic platform for assessing the performance of advanced ab initio methods.
Our study systematically evaluates a range of computational methods, including DFT, DFT+$U$, and the highly accurate but computationally expensive coupled-cluster method with single and double excitations (CCSD). We analyze the structural stability, electronic properties, and magnetic ordering of these 1D-TMOs, highlighting the strengths and limitations of each approach. In addition, we examine critical issues such as the influence of pseudopotentials vs full-potential calculations, convergence challenges, and the effects of structural distortions on electronic behavior.
In addition to their computational advantages, one-dimensional TMOs offer a useful platform for comparison with bulk systems. In summary, this paper provides a comprehensive analysis of ab initio methods applied to 1D-TMOs, offering insights into the challenges of modeling these systems. Prior studies have examined the magnetic and electronic properties of bulk 3d TMOs using DFT+$U$ and related methods Kuklin and Karttunen (2018); Qu et al. (2022); Trimarchi et al. (2018); Schrön et al. (2012); Gao et al. (2020). These works highlight the sensitivity of DFT+$U$ predictions to the choice of $U$ and the complexity of modeling strongly correlated systems. By comparing our 1D results with these bulk benchmarks, we assess which trends are preserved across dimensionalities and which are more sensitive to the structural environment. These comparisons are discussed in Section. III to further evaluate the transferability of ab initio methods. The paper is organized as follows: Section, II details the computational methods, Section. III presents our findings, including an analysis of convergence challenges and a comparison of DFT+$U$ with CCSD, and Section. IV concludes with a summary of our results.
@@ II COMPUTATIONAL METHODS
We use three computational codes for our calculations: Quantum ESPRESSO (QE) Giannozzi et al. (2009), a plane-wave pseudopotential code; PySCF Sun et al. (2018), a Python-based quantum chemistry framework; and FHI-aims Blum et al. (2009), an all-electron, full-potential electronic structure code. For density functional theory (DFT) calculations, we employ the Perdew–Burke–Ernzerhof (PBE) functional Perdew et al. (1996), a generalized gradient approximation (GGA) for the exchange-correlation energy. In DFT+$U$ calculations, we incorporate Dudarev’s formulation Dudarev et al. (1998) as implemented in QE.
@@
Our study focuses on first-row transition metal mono-oxides (TMOs, with TM = V, Cr, Mn, Fe, Co, and Ni) arranged in a one-dimensional (1D) chain structure along the $x$-direction (Fig. 1). Each chain is investigated in two magnetic configurations: ferromagnetic (FM) and antiferromagnetic (AFM). For AFM states, we use a minimal unit cell containing two formula units (four atoms) to account for magnetic ordering, and for the FM states, we adopt the same geometry unless explicitly stated. To minimize interactions between periodic images, we introduce a vacuum thickness of 30 atomic units (a.u.), and the Brillouin zone is sampled using a $4\times 1\times 1$ $k$-point mesh.
In QE, GBRV ultra-soft pseudopotentials are used Garrity et al. (2014) with a kinetic energy cutoff of 60 Ry for the plane-wave expansion. For all-electron calculations in FHI-aims, we select a tight-tier2 basis set. In PySCF calculations, we employ the Goedecker–Teter–Hutter (GTH) pseudopotential with the corresponding GTH-DZVP-MOLOPT-SR basis set (abbreviated as DZVP) VandeVondele and Hutter (2007); Gao et al. (2020), obtained from the CP2K library VandeVondele et al. (2005), to represent the valence electrons. Unrestricted CCSD calculations are carried out with initial reference states generated from unrestricted Hartree-Fock (UHF).
@@
Figure 1: The crystal structure of a one-dimensional (1D) transition-metal monoxide (TMO) chain, oriented along the $x$-axis, is shown. The transition metal (TM = V, Cr, Mn, Fe, Co, Ni) and oxygen atoms are represented by blue and red circles, respectively. The arrows indicate the magnetic moments of the transition metal atoms, illustrating two magnetic configurations: FM and AFM states.
@@ III.1 Instability and convergence issues
With the exception of the MnO chain, which shows stable convergence, all PBE and DFT+$U$ calculations—regardless of the DFT code used (i.e., PySCF, QE, and FHI-aims)—face significant wavefunction instability issues, often causing the self-consistent field (SCF) calculations to converge to an excited state instead of the ground state. Detecting such instabilities only requires comparing the total energy of the 1D-TMOs at different lattice parameters. Fig. 2(a) illustrates the variation in total energy with respect to the lattice parameter (TM–TM distance) for the CoO chain in the AFM state. Although the calculations converge successfully for all lattice parameters, some converge to unstable states, resulting in a zigzag energy plot rather than the expected smooth curve. This behavior indicates that some SCF calculations fail to reach a ground-state solution.
Projected density of states (PDOS) calculations reveal differences in $d$-orbital occupations between unstable and stable points, suggesting that these instabilities may arise from the limitations of ab initio methods in determining optimal electronic $d$-orbital occupations. Fig. 3 presents the spin-down charge density profile of CoO in a (011) surface cut, showing significant differences between stable and unstable configurations. This suggests that the $d$-orbital degrees of freedom in one-dimensional transition metal oxides (1D-TMOs) can contribute to instability in ab initio calculations.
@@
To overcome this issue, we adopt a straightforward yet effective stabilization procedure. We first examine the total energy profile [Fig. 2(a)] along with the magnetic moments of the transition metal atoms. When the energy profile shows irregularities—such as zigzag patterns or unexpected jumps—we identify lattice points that exhibit physically consistent solutions: smooth energy variation, lower total energy, symmetric magnetic moments on equivalent atoms, and correct magnetic ordering (e.g., AFM or FM as expected). These configurations are considered stable and used as reference points.
For unstable points (i.e., those with irregular energies or asymmetric local moments), we restart the SCF using the converged wavefunction from a nearby stable configuration as the initial guess. This targeted initialization helps the SCF escape from local excited-state minima and converge toward the correct ground-state solution. This approach is conceptually similar to Direct Inversion in the Iterative Subspace (DIIS) because it uses previously obtained stable solutions to guide convergence. Using this method, we successfully obtain smooth and physically meaningful total energy curves, as shown in Fig. 2(b). We apply this approach consistently across all DFT and HF calculations.
Figure 2: Variation in the total energy vs lattice constant (d) for CoO using the PBE method calculated by Quantum EEPRESSO, in the AFM phase :(a) Initial calculations showing instability in the wavefunction. (b) Final calculations after applying our trick to address the instability, demonstrating improved convergence.
Figure 3: (Colour online) The isosurface of the spin density for CoO in the FM state: (a) The spin density is isotropic for the unstable configuration. (b) An anisotropic spin density is observed for the stable configuration. (c) Shows the difference in spin density (subtracted). A slice of the spin density centered on the Co atom in the (011) surface is shown in the right panel.
@@ III.2 Structural properties
To study the electronic structure and magnetic properties of 1D-TMOs, we first determine their most stable atomic configurations. We perform a series of calculations for different lattice constants, considering both FM and AFM states. Fig. 4 illustrates how the total energy varies with the lattice parameter for each chain, computed using different ab initio methods.
Our results indicate that the AFM state is the most stable configuration for all chains, except CrO, when using both PBE and DFT+$U$. We obtain the optimized lattice constants, corresponding to the minimum energy values, by fitting the total energy data to the Birch–Murnaghan equation of state Murnaghan (1944); Birch (1947). These optimized lattice constants are then used to determine other properties, such as the magnetic moments of the 1D-TMOs.
To calculate the magnetization, we analyze the PDOS of these systems. Table. 1 summarizes the ground-state lattice constants, magnetization, and energy parameter ($\Delta E$) for each chain. The table clearly shows that the FM state generally results in slightly larger lattice constants and higher magnetic moments. However, within the PBE approach, the NiO chain is an exception, where the AFM state shows a larger lattice constant and higher magnetic moment compared to the FM state.
In addition, the magnetization of NiO is lower than expected within PBE, which can be attributed to its electronic structure near the Fermi level, where the orbitals are half-filled. This finding highlights the necessity of applying a Hubbard $U$ correction through DFT+$U$ calculations.
Figure 4: Variation of total energy with respect to the lattice constant (d). The reference energy ($E_{\mathrm{ref}}$) corresponds to the minimum energy of the stable phase. The calculations were performed using the PBE and DFT+$U$ methods with Quantum ESPRESSO and the CCSD method with PySCF for the FM and AFM phases of 1D-TMO chains.
Table 1: The calculated lattice parameters (d) and magnetic moments per ion in the AFM and FM phases, as well as the energy differences ($\Delta E$) between the two magnetic phases using the PBE method in QE, FHI-aims, and PySCF; the DFT+$U$ method in QE; and the CCSD method in PySCF. The absence of results for VO in PBE (PySCF) is due to the need for basis sets larger than TZVP to achieve convergence. For VO, we use fixed $U$ values of 4.8 eV, while for the other compounds, we employ a self-consistent $U$ parameter determined using DFPT.
@@
In the next step, to validate the pseudopotentials, we conducted a comparison between results obtained using the full-potential FHI-aims and pseudopotentials from QE, employing the GGA exchange-correlation functional. The calculated parameters are presented in Table. 1. The lattice parameters calculated using QE and FHI-aims show good agreement, with differences generally within 0.01–0.02 atomic units (a.u.), indicating the reliability of the present pseudopotentials approach in predicting the lattice constants. The Mulliken population analysis to estimate atomic spins also shows consistency between the two methods, though some differences are notable. For example, the magnetic moments of VO, CrO and MnO show small variations. The differences are more pronounced in compounds such as FeO (AFM), CoO (FM), and NiO (AFM). For the AFM state of FeO, the magnetization difference calculated by the two codes is 0.24 $\mu_{B}$; for the FM state of CoO it is 0.98 $\mu_{B}$, and for the AFM state of NiO the difference is about 0.2 $\mu_{B}$.
The energy differences between the AFM and FM phases ($\Delta E$) are generally consistent between the two methods, although there are some discrepancies. For VO, CrO, MnO, CoO, and NiO, the difference in $\Delta E$ between the methods is less than or equal to 0.05 eV in all cases. Only for FeO does this difference exceed 0.16 eV.
In summary, the GBRV pseudopotentials in QE generally provide results in close agreement with the full-potential FHI-aims calculations for lattice parameters, magnetic moments, and energy differences, validating their use in predicting the properties of these transition metal monoxide chains. The slight variations observed underscore the nuanced differences between pseudopotentials and all-electron methods, emphasizing the importance of method selection based on the specific requirements of a study.